One solution is 10% acid, another solution is 30% acid. find the volume of each solution that must be used to get a final solution of 40 liters that is 18% acid.
Answer by Maths68(1474)

Solution A
Amount = x ?
Concentration = 10% = 0.1
=======================================
Solution B
Amount =40- x ?
Concentration =30% = 0.3
=======================================
Resultant Solution
Amount = 40 liters
Concentration = 18%=0.18
=======================================
[Amount Solution A * Concentration A] + [Amount Solution B * Concentration of B] = Amount of Resultant * Concentration of resultant
========================================
Put the value in above equation
(x*0.1)+((40-x)*0.3=40*0.18
0.1x+12-0.3x=7.2
0.1x+12-0.3x=7.2
-0.2x=7.2-12
-0.2x=-4.8
Divide by (-0.2) both sides
-0.2x/-0.2=-4.8/-0.2
x=24 liters
=================================
Amount of solution A = 24 liters
Amount of Sultion B = 40-24=16 liters